Bwindi Impenetrable Forest (07.01.2016 to 09.30.2016)

You are here

PDF versionPDF version
Year: 
2016
Quarter: 
July - September
Communications activities: 
  • Edmond Kagmugisha, a Msc student from Mbarara University of Science and Technology, has started his data analysis using TEAM vegetation data. He is supervised by Associate Professor Robert Bitariho, the director of ITFC. The research title is: Estimating the above ground biomass and carbon stocks in Bwindi Impenetrable National Park.
  • Emmanuel Akampulira, the TEAM site manager in Bwindi, also intends to use the TEAM camera pictures for his Msc research. He intends to assess the relationship between the distribution of animals and illegal activities in Bwindi.

 

 

Unusual Events at the TEAM Site: 

Camera trapping pictures in Array two captured elephants in the northern sector. This occurrence has not been reported in Bwindi in more than 20 years. Elephants in Bwindi have primarily been recorded in the southern sector in the past 25 years. 

New Species: 

None

Protocol Activities: 

Vegetation sampling

Data from the vegetation plots were entered and uploaded on the TEAM portal in the previous quarters.  

Climate monitoring protocol

All climate data and maintenance logs were uploaded on to the TEAM portal. Bi-weekly maintenance of the climate station has been carried out in accordance with the protocol. We received two Temp/RH sensors from Campbell Scientific and they are both working well. Unfortunately the third sensor is malfunctioning and cannot be replaced by Campbell scientific because it’s no longer under warranty.  We plan to purchase another sensor in the next project year.

Terrestrial vertebrate monitoring protocol

Camera trapping was completed. We set 30 camera traps for 30 days in array one and 30 camera traps for 30 days in array two, giving a total of 60 camera traps set for 60 days.  Images from both arrays were downloaded into DeskTEAM computer program and processed.  The tpk files were copied onto two SD cards and posted to Conservation International offices in Washington.  

Protocol Problems: 
  • Since the beginning of the year, we have had malfunctioning Temp/RH sensors. We received two Temp/RH sensors from Campbell Scientific to replace the faulty ones and they are both working well. Unfortunately the third sensor is now malfunctioning and cannot be replaced by Campbell scientific because it’s no longer under warranty.  We plan to purchase another sensor in the next project year. This problem was communicated to the TEAM Technical Director.
  • Camera trap 12 was approached from behind, unlocked, and the wires connecting the battery compartment to camera was cut. This was during deployment in array 2 and we do not know who did this.
Schedule Problems: 

None

Logistical Problems: 
  • Currently we are in need of camping tents for our fieldwork teams. The tents available are old and torn and we cannot continue to use them for TEAM project.
  • We also have very old backpacks that are not helpful in the field. 
  • Our GPS sets have also grown old and we currently have one GPS that functions well. 
Suggestions: 
  • For the equipment listed above, we suggest that we use some of the money in the next project budget to buy them. 
  • It's also our request that funds for the next year be transferred early to allow timely and easy execution of activities.